# On the Entropy Calibration of Language Models

### v1

@@ Abstract
We study the problem of entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text. Past work found that models are miscalibrated, with entropy per step increasing (and text quality decreasing) as generations grow longer. This error accumulation is a fundamental problem in autoregressive models, and the standard solution is to truncate the distribution, which improves text quality at the cost of diversity. In this paper, we ask: is miscalibration likely to improve with scale, and is it theoretically possible to calibrate without tradeoffs? To build intuition, we first study a simplified theoretical setting to characterize the scaling behavior of miscalibration with respect to dataset size. We find that the scaling behavior depends on the power law exponent of the data distribution — in particular, for a power law exponent close to 1, the scaling exponent is close to 0, meaning that miscalibration improves very slowly with scale. Next, we measure miscalibration empirically in language models ranging from 0.5B to 70B parameters. We find that the observed scaling behavior is similar to what is predicted by the simplified setting: our fitted scaling exponents for text are close to 0, meaning that larger models accumulate error at a similar rate as smaller ones. This scaling (or, lack thereof) provides one explanation for why we sample from larger models with similar amounts of truncation as smaller models, even though the larger models are of higher quality. However, truncation is not a satisfying solution because it comes at the cost of increased log loss. In theory, is it even possible to reduce entropy while preserving log loss? We prove that it is possible, if we assume access to a black box which can fit models to predict the future entropy of text.
@@ 1 Introduction
We study entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text. This definition is a natural notion of calibration for generative tasks, and is much more challenging than calibration for classification tasks because the output space is exponentially large. While we will discuss this definition in more detail later, one basic requirement for calibration is that the model’s entropy per step should be stable over the generation.https://github.com/stevenxcao/entropy-calibration
Unfortunately, autoregressive language models are not stable. Entropy calibration was first studied by braverman2020calibration, who found that language models have entropy per step increasing as generations grow longer. This entropy blowup is accompanied by an increase in generation errors, as is also observed in basu2021mirostatneuraltextdecoding. While entropy calibration specifically is not well studied, generation instability more broadly has been the subject of many papers (williams1989learning; Ranzato2015SequenceLT; Welleck2019NeuralTG). From this line of work has emerged a suite of distribution truncation techniques, which have become standard practice in language model sampling (fan-etal-2018-hierarchical; holtzman2020curious; hewitt-etal-2022-truncation). These techniques suppress low probability tokens to improve quality at the cost of diversity (hashimoto-etal-2019-unifying; zhang-etal-2021-trading).
It is not fully satisfying that to make generation stable, we must sacrifice diversity. Diversity is especially important for difficult tasks where we must aggregate multiple answers (wang2024mixtureofagentsenhanceslargelanguage; brown2024largelanguagemonkeysscaling), as well as for synthetic data generation, which has seen a resurgence of interest as the community has begun worrying about running out of internet data (wang-etal-2023-self-instruct; gunasekar2023textbooksneed; maini-etal-2024-rephrasing). Therefore, it is natural to ask: do we expect generation stability to improve with scale? If not, is it at least theoretically possible to calibrate without sacrificing diversity?
To build intuition, we first study a simplified theoretical setting, where instability comes from the fact that the model might generate a token that it saw only a few times during training. This unfamiliar token then derails subsequent steps when it is fed back into the context autoregressively. Drawing on classic results, we calculate a scaling exponent capturing how quickly the probability of generating a rare token decreases with the number of training examples (good-1953-population; karlin-1967-central). We find that this exponent depends on how heavy-tailed the data distribution is: in particular, for power law exponents close to 1, as is typical for human text (zipf-1936-psycho; zipf-1949-human), the scaling exponent is close to 0. Therefore, this setup predicts that stability in generation improves very slowly with scale.
@@
If even large models suffer from error accumulation, why are reasoning and instruction-tuned models able to produce long, coherent outputs? We find that much like distribution truncation, instruction tuning reduces entropy at the cost of increased log loss, with the largest models now having entropy too low. This tradeoff relates to past work which found that alignment degrades model capabilities, a phenomenon known as the alignment tax (ouyang2022training; bai2022helpful; lin-etal-2024-mitigating).
@@ 2 Preliminaries
We first review key definitions and properties for entropy calibration, introduced in braverman2020calibration. Our setup is as follows: we are given prompts $X\in\mathcal{X}$ drawn from some prompt distribution $X\sim q$, and responses $Y\in\mathcal{Y}$ drawn from the true conditional distribution $Y\sim p^{*}_{X}$. For example, $X$ might contain a description of a coding task, while $Y$ contains a solution to the task. We then train a language model $\hat{p}:\mathcal{X}\to\Delta^{\mathcal{Y}}$ to fit the true conditional distribution $p^{*}$. We say that $\hat{p}$ is entropy calibrated if its entropy over generations is equal to its log loss:
@@
To build intuition for this definition, entropy can be thought of as a measure of the model’s uncertainty, which should be calibrated to match the actual loss it incurs on real data. This definition mirrors that of binary calibration, and we derive this connection more formally in Appendix B. Qualitatively, if a model is underconfident, then its generations have too much entropy and appear incoherent; if it is overconfident, then its generations have too little entropy and appear repetitive (braverman2020calibration; basu2021mirostatneuraltextdecoding); see Appendix E for a replication of this finding and Appendix D for examples. Entropy calibration is then the problem of adjusting the entropy to be just right. Empirically, braverman2020calibration found that neural autoregressive language models have entropy too high: entropy per step matches the log loss at earlier steps but increases as the generation grows.
@@
Why does entropy per step grow with the length of the generation? The main problem, as has been observed in empirical work, is that autoregressive language models accumulate error during generation. At training time, models are given input from the true distribution and asked to produce only a single additional token. In contrast, models must generate multiple tokens at deployment time, which they do by producing one token at a time and taking their own production as subsequent input. Therefore, even models with very low single-step error can degrade over multiple steps as they take their own slightly erroneous outputs as input and accumulate errors (see, e.g., Ranzato2015SequenceLT, Welleck2019NeuralTG, holtzman2020curious for error accumulation in language modeling; and daume2009search, ross2010efficient, ross2011reduction for imitation learning). This intuition is formalized in the context of entropy calibration in Corollary 4.2 of braverman2020calibration, which states that for a model with $\varepsilon$ KL divergence to the true distribution, the entropy at step $t$ can deviate as much as $\varepsilon+\sqrt{\varepsilon t}$ from the log loss, growing with $t$.
@@
How does one calibrate the entropy? Unlike binary and multiclass calibration, entropy calibration is challenging because the models have an exponentially large output space. In practice, practitioners use a number of distribution truncation methods, each of which uses a different heuristic to suppress low probability tokens in each generation step. Some standard methods include temperature reduction, top-k sampling (fan-etal-2018-hierarchical), top-p sampling  (holtzman2020curious), and min-p sampling (hewitt-etal-2022-truncation). These methods improve text quality at the cost of diversity (hashimoto-etal-2019-unifying; zhang-etal-2021-trading; pillutla2021mauve; welleck2024decodingmetagenerationinferencetimealgorithms). Following hashimoto-etal-2019-unifying, we define a model’s diversity to be its log loss on reference documents. The intuition behind this definition is that log loss (also known as cross entropy or forward KL) is a coverage metric: to attain low log loss, the model must “cover” as much as the reference distribution as possible. Our goal, then, is to calibrate entropy to match log loss without also causing the log loss to also increase.
@@
In theory, braverman2020calibration show that one can calibrate entropy while preserving log loss via globally normalized temperature scaling, where the adjusted model is given by $\hat{p}_{\tau}(y_{1},...,y_{L})\propto\hat{p}(y_{1},...,y_{L})^{1/\tau}$. Unfortunately, this adjustment is intractable to compute because it involves normalizing over the entire output space. It remains unclear, then, whether this goal is possible in polynomial time. Specifically, we wish to take in a model $\hat{p}$ and produce a calibrated model $\tilde{p}$ with at most $\varepsilon$ entropy calibration error per step, as well as log loss at most that of the original model $\hat{p}$:
@@ 3 Intuition: Singleton Mass in Power Law Data
Before putting in the work to develop better calibration algorithms, it is natural to first ask whether we expect miscalibration to automatically improve with scale, as we train larger models on more data. To gain intuition, we first explore this question in a simplified theoretical setting. We define the setup to capture the following hypothesis regarding error accumulation (see, e.g., hewitt-etal-2022-truncation): because the language distribution is heavy-tailed, the model must assign non-zero probability to a large number of rare tokens when fitting the data. However, if it happens to generate one such rare token, the model derails when that token is fed back into the context autoregressively, leading to a jump in entropy. Over many generation steps, then, the model will eventually derail. The degree of instability then depends on the probability of producing a rare token.
@@
In this setting, the expected entropy per step grows with slope proportional to the probability of emitting a rare token (see Appendix B). Computing the rare token mass in power law data is a classic problem, and we can compute the asymptotic scaling exponent with respect to the number of training examples $m$ as follows (good-1953-population; karlin-1967-central):
@@
We provide the derivation in the appendix. The key takeaway from this proposition is that the derailing probability scales as $m^{1/\alpha-1}$, which is very slow if the power law exponent $\alpha$ is close to $1$, as is typical for text (zipf-1936-psycho; zipf-1949-human). The reason for this slow scaling is that as $m$ increases, there are always more rare items to be sampled from the tail of the distribution. In practice, of course, we are not training unigram models, but the same intuition holds if we posit that semantic concepts in text are similarly heavy tailed: as larger models are trained on more data, there will always be new rare phenomena that they see during training only once. These phenomena are then memorized, and potentially generated at deployment, derailing the model.
While asymptotic analysis gives us a clean expression, we can also estimate the scaling exponent in simulation for finite values of $m$ and $v$ (see Figure 1). We find that the actual slopes are close to the asymptotic expression, up to $m\approx v/3$. We also calculate the power law exponent for our three datasets, finding that it is around 1 for WikiText and WritingPrompts and $1.5$ for CodeContests, which predicts slow scaling for the first two datasets and slow-to-moderate scaling for the third.
@@ 4 Experiments: Miscalibration in Large Language Models
Next, we measure miscalibration empirically in large language models. We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (qwen2025qwen25technicalreport), Llama 3 (1B, 3B, 8B, 70B) (grattafiori2024llama3herdmodels), Llama 2 (7B, 13B, 70B) (touvron2023llama2openfoundation), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (biderman2023pythiasuiteanalyzinglarge)) applied to the three datasets listed below. In each setting, we use 5000 examples and limit samples to 1024 tokens; see Appendix C for more experimental details. We primarily study base models because we are interested in the problem of modeling human text; we study the effect of instruction tuning in Section 4.3.
@@
    WikiText-103 (merity2016pointersentinelmixturemodels): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
@@
    WritingPrompts (fan-etal-2018-hierarchical): given a prompt from r/writingprompts along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
@@
    CodeContests (li-2022-competition): given a coding problem from one of five websites and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
@@ 4.1 Miscalibration scaling in base models
Past work has found that many model capabilities improve predictably with model size, with task loss and model size following a linear relationship when plotted on a log scale (kaplan2020scalinglawsneurallanguage; hoffmann2022trainingcomputeoptimallargelanguage). We use a similar methodology to study the relationship between entropy miscalibration and model size. If model size and dataset size are scaled proportionally, Proposition 3.1 suggests a scaling law of $\log\text{EntCE}=(1/\alpha-1)\log m+C$, where $\alpha$ is the power law exponent of the data distribution and $m$ is the parameter count. Does the actual data also follow a clean scaling law, and how close is the scaling exponent to that predicted by the simplified setting?
@@ 4.2 Entropy over time
First, we find that for each model and dataset, the log loss is mostly constant or slightly decreasing over time. Past papers use a model’s log loss to estimate the actual entropy of the underlying text, as the former is an upper bound for the latter that grows tighter if the model is more accurate. This part of the plot replicates past findings that the entropy per step of human text is stable over time, also known as the entropy rate constancy principle (genzel-charniak-2002-entropy; verma-etal-2023-revisiting).
@@ 4.3 Calibration-diversity tradeoffs
First, we find that reducing temperature below $1$ reduces entropy but increases log loss, replicating similar findings in past work (hashimoto-etal-2019-unifying). Furthermore, the temperature attaining zero calibration error is similar across model sizes, which makes sense given that they are similarly miscalibrated. We find that instruction tuning also reduces entropy while increasing log loss, which is consistent with past work showing that instruction tuning harms diversity (ghosh2024closer). Unlike temperature scaling, the magnitude of the effect varies across model sizes, with larger models experiencing both a larger reduction in entropy and larger increase in log loss. However, this pattern is not robust across model families (see Appendix E). Further work with more controlled instruction tuning would be necessary to explore this relationship further. These experiments reconcile our previous finding, that even large models accumulate errors, with the fact that in practice, one can use truncation or post-training to generate long, coherent pieces of text. The tradeoff is that each of these mitigations comes at the cost of diversity.

### v2

@@ Abstract
We study the problem of entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text. Past work found that models are miscalibrated, with entropy per step increasing as generations grow longer, due to error accumulation. To calibrate the model and improve text quality, it has become standard practice to truncate the distribution, but this approach reduces output diversity, which we would like to avoid. Therefore, in this paper, we ask: does miscalibration improve automatically with scale, and if not, is it theoretically possible to calibrate without tradeoffs? To build intuition, we first study a simplified theoretical setting to characterize the scaling behavior of miscalibration with respect to dataset size. We find that the rate of scaling depends on the power law exponent of the data distribution — in particular, for a power law exponent close to 1, the scaling exponent is close to 0, meaning that miscalibration improves very slowly with scale. Next, we measure miscalibration empirically in language models ranging from 0.5B to 70B parameters. We find that the observed scaling behavior is similar to what is predicted theoretically: our fitted scaling exponents for text are close to 0, meaning that larger models accumulate error at a similar rate as smaller ones. This scaling (or, lack thereof) provides one explanation for why we sample from larger models with similar amounts of truncation as smaller models, even though the larger models are of higher quality. However, truncation is not a satisfying solution because it comes at the cost of increased log loss. In theory, is it even possible to reduce entropy while preserving log loss? We prove that it is possible, if we assume access to a black box which can fit models to predict the future entropy of text.
@@ 1 Introduction
We study entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text.https://github.com/stevenxcao/entropy-calibration This definition is a natural notion of calibration for generative tasks, and is more challenging than calibration for classification tasks because the output space is exponentially large. Braverman et al. (2020) were the first to study entropy calibration in language models, and they found that models are miscalibrated: entropy per step for model generations increases with the length of the document, in contrast with log loss on human text which is roughly flat over the length of the document (Genzel and Charniak, 2002; Verma et al., 2023). Entropy calibration can be thought of as a quantitative formalization of the well-known “error accumulation” or “teacher forcing” problem: entropy rises when the model generates erroneous tokens which are fed back into the context, derailing the generation (Williams and Zipser, 1989; Ranzato et al., 2016; Basu et al., 2021; Hewitt et al., 2022, also see Appendix E). Therefore, we study entropy calibration to gain fundamental insights into improving generation quality.
To correct error accumulation and calibrate the model, it has become standard practice to truncate the next token distribution (Fan et al., 2018; Holtzman et al., 2020; Hewitt et al., 2022), suppressing low probability tokens to improve quality at the cost of diversity (Hashimoto et al., 2019; Zhang et al., 2021). This solution is not satisfying: diversity is especially important for difficult tasks where we must aggregate multiple answers (Wang et al., 2024; Brown et al., 2024), as well as for synthetic data generation, which has seen a resurgence of interest as the community has begun worrying about running out of internet data (Wang et al., 2023; Gunasekar et al., 2023; Maini et al., 2024). Therefore, it is natural to ask: do we expect miscalibration to improve with scale? If not, is it at least theoretically possible to calibrate without sacrificing diversity?
To build intuition, we first study a simplified theoretical setting, where instability comes from the fact that the model might generate a token that it saw only a few times during training. This unfamiliar token then derails subsequent steps when it is fed back into the context autoregressively. Drawing on classic results, we calculate a scaling exponent capturing how quickly the probability of generating a rare token decreases with the number of training examples (Good, 1953; Karlin, 1967). We find that this exponent depends on how heavy-tailed the data distribution is: in particular, for power law exponents close to 1, as is typical for human text (Zipf, 1936, 1949), the scaling exponent is close to 0. Therefore, this setup predicts that stability in generation improves very slowly with scale.
@@
If even large models suffer from error accumulation, why are reasoning and instruction-tuned models able to produce long, coherent outputs? We find that much like distribution truncation, instruction tuning reduces entropy at the cost of increased log loss, with the largest models now having entropy too low. This tradeoff relates to past work which found that alignment degrades model capabilities, a phenomenon known as the alignment tax (Ouyang et al., 2022; Bai et al., 2022; Lin et al., 2024).
@@ 2 Preliminaries
We first review key definitions and properties for entropy calibration, introduced in Braverman et al. (2020). Our setup is as follows: we are given prompts $X\in\mathcal{X}$ drawn from some prompt distribution $X\sim q$, and responses $Y\in\mathcal{Y}$ drawn from the true conditional distribution $Y\sim p^{*}_{X}$. For example, $X$ might contain a description of a coding task, while $Y$ contains a solution to the task. We then train a language model $\hat{p}:\mathcal{X}\to\Delta^{\mathcal{Y}}$ to fit the true conditional distribution $p^{*}$. We say that $\hat{p}$ is entropy calibrated if its entropy over generations is equal to its log loss:
@@
To build intuition for this definition, entropy can be thought of as a measure of the model’s uncertainty, which should be calibrated to match the actual loss it incurs on real data. This definition mirrors that of binary calibration, and we derive this connection more formally in Appendix B. Qualitatively, if a model is underconfident, then its generations have too much entropy and appear incoherent; if it is overconfident, then its generations have too little entropy and appear repetitive (Braverman et al., 2020; Basu et al., 2021); see Appendix E for a replication of this finding and Appendix D for examples. Entropy calibration is then the problem of adjusting the entropy to be just right. Empirically, Braverman et al. (2020) found that neural autoregressive language models have entropy too high: entropy per step matches the log loss at earlier steps but increases as the generation grows.
@@
Why does entropy per step grow with the length of the generation? The main problem, as has been observed in empirical work, is that autoregressive language models accumulate error during generation. At training time, models are given input from the true distribution and asked to produce only a single additional token. In contrast, models must generate multiple tokens at deployment time, which they do by producing one token at a time and taking their own production as subsequent input. Therefore, even models with very low single-step error can degrade over multiple steps as they take their own slightly erroneous outputs as input and accumulate errors (see, e.g., Ranzato et al. (2016), Welleck et al. (2020), Holtzman et al. (2020) for error accumulation in language modeling; and Daumé et al. (2009), Ross and Bagnell (2010), Ross et al. (2011) for imitation learning). This intuition is formalized in the context of entropy calibration in Corollary 4.2 of Braverman et al. (2020), which states that for a model with $\varepsilon$ KL divergence to the true distribution, the entropy at step $t$ can deviate as much as $\varepsilon+\sqrt{\varepsilon t}$ from the log loss, growing with $t$.
@@
How does one calibrate the entropy? Unlike binary and multiclass calibration, entropy calibration is challenging because the models have an exponentially large output space. In practice, practitioners use a number of distribution truncation methods, each of which uses a different heuristic to suppress low probability tokens in each generation step. Some standard methods include temperature reduction, top-k sampling (Fan et al., 2018), top-p sampling  (Holtzman et al., 2020), and min-p sampling (Hewitt et al., 2022). These methods improve text quality at the cost of diversity (Hashimoto et al., 2019; Zhang et al., 2021; Pillutla et al., 2021; Welleck et al., 2024). Following Hashimoto et al. (2019), we define a model’s diversity to be its log loss on reference documents. The intuition behind this definition is that log loss (also known as cross entropy or forward KL) is a coverage metric: to attain low log loss, the model must “cover” as much as the reference distribution as possible. Our goal, then, is to calibrate entropy to match log loss without also causing the log loss to increase.
@@
In theory, Braverman et al. (2020) show that one can calibrate entropy while preserving log loss via globally normalized temperature scaling, where the adjusted model is given by $\hat{p}_{\tau}(y_{1},...,y_{L})\propto\hat{p}(y_{1},...,y_{L})^{1/\tau}$. Unfortunately, this adjustment is intractable to compute because it involves normalizing over the entire output space. It remains unclear, then, whether this goal is possible in polynomial time. Specifically, we wish to take in a model $\hat{p}$ and produce a calibrated model $\tilde{p}$ with at most $\varepsilon$ entropy calibration error per step, as well as log loss at most that of the original model $\hat{p}$:
@@ 3 Intuition: Singleton Mass in Power Law Data
Before putting in the work to develop better calibration algorithms, it is natural to first ask whether we expect miscalibration to automatically improve with scale, as we train larger models on more data. To gain intuition, we first explore this question in a simplified theoretical setting. We define the setup to capture the following hypothesis regarding error accumulation (see, e.g., Hewitt et al. (2022)): because the language distribution is heavy-tailed, the model must assign non-zero probability to a large number of rare tokens when fitting the data to achieve low log loss. However, if it happens to generate one such rare token, the model derails when that token is fed back into the context autoregressively, leading to a jump in entropy. Over many generation steps, then, the model will eventually derail. The degree of instability then depends on the probability of producing a rare token.
@@
In this setting, the expected entropy per step grows with slope proportional to the probability of emitting a rare token (see Appendix B). Computing the rare token mass in power law data is a classic problem, and we can compute the asymptotic scaling exponent with respect to the number of training examples $m$ as follows (Good, 1953; Karlin, 1967):
@@
We provide the derivation in the appendix. The key takeaway from this proposition is that the derailing probability scales as $m^{1/\alpha-1}$, which is very slow if the power law exponent $\alpha$ is close to $1$, as is typical for text (Zipf, 1936, 1949). The reason for this slow scaling is that as $m$ increases, there are always more rare items to be sampled from the tail of the distribution. In practice, of course, we are not training unigram models, but the same intuition holds if we posit that semantic concepts in text are similarly heavy tailed: as larger models are trained on more data, there will always be new rare phenomena that they see during training only once. These phenomena are then memorized, and potentially generated at deployment, derailing the model.
While asymptotic analysis gives us a clean expression, we can also estimate the scaling exponent in simulation for finite values of $m$ and $v$. We find that the non-asymptotic simulated slopes are close to the asymptotic expression as long is $m$ is smaller than $v/3$ (see Appendix E). We also calculate the power law exponent for our three datasets, finding that it is around 1 for WikiText and WritingPrompts and $1.5$ for CodeContests, which predicts slow scaling for the first two datasets and slow-to-moderate scaling for the third.
@@ 4 Experiments: Miscalibration in Large Language Models
Next, we measure miscalibration empirically in large language models. We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (Qwen et al., 2025), Llama 3 (1B, 3B, 8B, 70B) (Grattafiori et al., 2024), Llama 2 (7B, 13B, 70B) (Touvron et al., 2023), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (Biderman et al., 2023)) applied to the three datasets listed below. In each setting, we use 5000 examples and limit samples to 1024 tokens; see Appendix C for more experimental details. We primarily study base models because we are interested in the problem of modeling human text; we study the effect of instruction tuning in Section 4.3.
@@
    WikiText-103 (Merity et al., 2017): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
@@
    WritingPrompts (Fan et al., 2018): given a prompt from r/writingprompts along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
@@
    CodeContests (Li et al., 2022): given a coding problem from one of five websites and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
@@ 4.1 Miscalibration scaling in base models
Past work has found that many model capabilities improve predictably with model size, with task loss and model size following a linear relationship when plotted on a log scale (Kaplan et al., 2020; Hoffmann et al., 2022). We use a similar methodology to study the relationship between entropy miscalibration and model size. If model size and dataset size are scaled proportionally, Proposition 3.1 suggests a scaling law of $\log\text{EntCE}=(1/\alpha-1)\log m+C$, where $\alpha$ is the power law exponent of the data distribution and $m$ is the parameter count. Does the actual data also follow a clean scaling law, and how close is the scaling exponent to that predicted by the simplified setting?
@@ 4.2 Entropy over time
First, we find that for each model and dataset, the log loss is mostly constant or slightly decreasing over time. Past papers use a model’s log loss to estimate the actual entropy of the underlying text, as the former is an upper bound for the latter that grows tighter if the model is more accurate. This part of the plot replicates past findings that the entropy per step of human text is stable over time, also known as the entropy rate constancy principle (Genzel and Charniak, 2002; Verma et al., 2023).
@@ 4.3 Calibration-diversity tradeoffs
First, we find that reducing temperature below $1$ reduces entropy but increases log loss, replicating similar findings in past work (Hashimoto et al., 2019). Furthermore, the temperature attaining zero calibration error is similar across model sizes, which makes sense given that they are similarly miscalibrated. We find that instruction tuning also reduces entropy while increasing log loss, which is consistent with past work showing that instruction tuning harms diversity (Ghosh et al., 2024). Unlike temperature scaling, the magnitude of the effect varies across model sizes, with larger models experiencing both a larger reduction in entropy and larger increase in log loss. However, this pattern is not robust across model families (see Appendix E). Further work with more controlled instruction tuning would be necessary to explore this relationship further. These experiments reconcile our previous finding, that even large models accumulate errors, with the fact that in practice, one can use truncation or post-training to generate long, coherent pieces of text. The tradeoff is that each of these mitigations comes at the cost of diversity.